OpenStudy (anonymous):

PLEASE help! Will medal and become a fan! check my answer! 2. Which of the following quartic functions has x=-2 and x=-3 as its only two real zeros? A x^4-5x^3+7x^2-5x+6 B x^4+5x^3+7x^2+5x+6 ☺ C x^4+5x^3+7x^2-5x-6 D x^4-5x^3+7x^2+5x-

OpenStudy (anonymous):

If a polynomial has those two zeros, you know it must factorise like this \[(x+2)(x+3)\times something\] So you have \[(x+2)(x+3) = x^2 + 5x + 6\] You want something with an x^4 so why not try multiplying this by x^2? \[x^2(x+2)(x+3) = x^4 + 5x^3 + 6x^2\] Now already we can eliminate two answers! A and D both have \[x^4 - 5x^3 ...\] so they are wrong. So we're left with B or C We need a 5x and a 6 in the answer, but it's either plus or minus 1 looking at B or C (because (x+2)(x+3) gives you a 5x and a 6, right?). So let's look at the x^2 term. It's 7x^2 for both of them and at the moment we have 6x^2 term. Now you know it has to be \[(x^2+1)(x+2)(x+3) = x^4 + 5x^3 + 7x^2 + 5x + 6\] and that's B